OpenStudy (anonymous):

determine whether the functions even, odd or neither. f(x)= x^2sinX. can any one explain how to check if this is even or odd pls.

OpenStudy (anonymous):

Good question. x^2 is even, it is symmetric around x = 0 and so (-1)^2 = 1^2, for example sin x is odd, it is not symmetric, but sin x = - sin (-x) OK so far?

OpenStudy (anonymous):

put -x (-x)^2sin(-X) = -x^2sinX sign changes so odd otherwise even

OpenStudy (anonymous):

Yes. f(-x) = - f(x), so x^2 sin x is odd.

OpenStudy (anonymous):

got it thank u!

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

is this right? f(x)= x^3+cosX f(-x)=(-x)^3+cos(-x)-------->(-X^3)+(-cosX)----->so f(-x)= -X^3-cosX thus this is odd because the sing changes??

OpenStudy (anonymous):

absolutely correct :)

OpenStudy (anonymous):

why could the answer be neither?

OpenStudy (anonymous):

oops, my badd f(-x)=(-x)^3+cos(-x)-------->(-X^3)+(+cosX)----->so f(-x)= -X^3+cosX neither

OpenStudy (anonymous):

gtg

OpenStudy (anonymous):

Think about what happens to f(x)= x^3+cosX for negative x

OpenStudy (anonymous):

take x = pi/4 f(pi/4) = (pi/4)^3 + 1/sqrt(2) f(-pi/4) = -(pi/4)^3 + 1/sqrt(2) so f(x) is not equal to f(-x) but neither is it equal to -f(-x)

OpenStudy (anonymous):

I dont understand how we get a +x when we cube a -x or how we find the x from -x...

OpenStudy (anonymous):

(-x)^3 = (-1)^3 x^3 = -1 x^3 = -x^3

OpenStudy (anonymous):

make that last -(x^3)

OpenStudy (anonymous):

hum I mean is see -(x^3) the same as -x^3...Im sorry, I don't get it. =[

OpenStudy (anonymous):

An even function is a function where f(-x) = f(x). So x^2 is an even function. OK?

OpenStudy (anonymous):

And an odd function is a function where f(-x) = -f(x). But some functions are neither even nor odd. OK?

OpenStudy (anonymous):

Try a product like f(x) = x^2 cos x (-x)^2 = x^2 cos x = cos -x So f(-x) = (-x)^2 cos -x = x^2 cos x = f(x) so it's even. OK?

OpenStudy (anonymous):

Lost you I guess.

OpenStudy (anonymous):

I understand how the powers work but like on cos x= cos -x we are changing it to f(-x) which I see but I don't see how it ends up being positive--cosx

OpenStudy (anonymous):

Umm

OpenStudy (anonymous):

If you look at the figure for the cosine function, do you see how it is symmetrical across the y axis?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

That symmetry is exactly the same as saying that cos(x) = cos(-x)

OpenStudy (anonymous):

Or look at this figure, I just grabbed it quickly so it's not properly labeled but still.

OpenStudy (anonymous):

A negative angle h is a rotation below the x-axis. But the cosine is still positive.

OpenStudy (anonymous):

ok I see, have a better idea now. thank you!

OpenStudy (anonymous):

yw